jakes surf shop rents surfboards for $6.00 plus $3.00 per hour. ritas surf shop rents them for $9.00 plus $2.50 per hour. which is cheaper for which time intervals?
Answer by stanbon(75887)

jakes surf shop rents surfboards for $6.00 plus $3.00 per hour. ritas surf shop rents them for $9.00 plus $2.50 per hour. which is cheaper for which time intervals?
Let "x" be the number of hours of rental.
Jakes Charges:
CJ= 6+3x
Ritas Charges:
CR=9+2.5x
Their charges are the same when 6+3x=9+2.5x
0.5x=3
x=6 hrs.
Consider CJ(4)=6+3(4)=$18
and CR(4)=9+2.5(4)=$19
Looks like Jake is cheaper for x hrs <6
Cheers,
Stan H.